OpenStudy (logan13):

Given quadrilateral LMNO ~ quadrilateral RSTU and LM = 8, NO = 12, and RS = 7, what is TU?

OpenStudy (logan13):

@jhonyy9

OpenStudy (logan13):

@rebeccaxhawaii

jhonyy9 (jhonyy9):

write the proportionalitie of coresponding sides do you know it how ?

OpenStudy (logan13):

no

jhonyy9 (jhonyy9):

to LM corespond LS so how will be than ?

OpenStudy (logan13):

idk

jhonyy9 (jhonyy9):

LM NO --- = --- RS TU

OpenStudy (logan13):

so i multiply numbers

OpenStudy (logan13):

*letters

jhonyy9 (jhonyy9):

substiitute these given values and calculi the TU

jhonyy9 (jhonyy9):

8/7 = 12/x x = ?

OpenStudy (logan13):

1.14285714286

OpenStudy (mathstudent55):